OpenStudy (anonymous):

simplify the rational expression. state any excluded values. x-2 over x^2+3x-10

OpenStudy (anonymous):

wait so it is x-2/x^2 + 3x+10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

factor x^2+3x+10 first

OpenStudy (anonymous):

no its 3x-10 not plus

OpenStudy (raden):

factor out the denominator : x^2+3x-10 = (x+5)(x-2) now, u have (x-2)/(x^2+3x-10) = (x-2)/(x+5)(x-2) then cancel out the (x-2), the rest is ur answer

OpenStudy (anonymous):

thank you

OpenStudy (raden):

yw